Given the points (2,6) and (-4,10) found on line AB; find the slope-intercept form of this line AB.
Answer by stanbon(75887)

Given the points (2,6) and (-4,10) found on line AB; find the slope-intercept form of this line AB.
----------------------
slope = (10-6)/(-4-2) = 4/-6 = -2/3
-----
Form:: y = mx + b
Solve for "b"::
6 = (-2/3)2 + b
6 = -4/3 + b
b = 22/3
-----
Equation:
y = (-2/3)x + (22/3)
-----------------------
Cheers,
Stan H.
============